# Instruction-Tuned LLMs Succeed in Document-Level MT Without Fine-Tuning—But BLEU Turns a Blind Eye

### v1

@@ Abstract
Large language models (LLMs) have excelled in various NLP tasks, including machine translation (MT), yet most studies focus on sentence-level translation. This work investigates the inherent capability of instruction-tuned LLMs for document-level translation (docMT). Unlike prior approaches that require specialized techniques, we evaluate LLMs by directly prompting them to translate entire documents in a single pass. Our results show that this method improves translation quality compared to translating sentences separately, even without document-level fine-tuning. However, this advantage is not reflected in BLEU scores, which often favor sentence-based translations. We propose using the LLM-as-a-judge paradigm for evaluation, where GPT-4 is used to assess document coherence, accuracy, and fluency in a more nuanced way than n-gram-based metrics. Overall, our work demonstrates that instruction-tuned LLMs can effectively leverage document context for translation. However, we caution against using BLEU scores for evaluating docMT, as they often provide misleading outcomes, failing to capture the quality of document-level translation.  [^1] Our code and the outputs from GPT4-as-a-judge will be released upon acceptance.
Instruction-Tuned LLMs Succeed in Document-Level MT Without Fine-Tuning—But BLEU Turns a Blind Eye
@@ 1 Introduction
A key challenge in our research is the evaluation of document-level machine translation (docMT). Traditional metrics like BLEU, ChrF, and TER Papineni et al. (2002); Popović (2015); Snover et al. (2006), though widely used, often poorly correlate with human judgment Freitag et al. (2022), especially in docMT, where maintaining coherence and logical flow across a document is essential—something n-gram overlap struggles to capture. Metrics like cTT, aZPT, and Blonde Jiang et al. (2021); Wang et al. (2023) address specific aspects such as terminology consistency and zero-pronoun accuracy, but still rely heavily on word matching and symbolic statistics. We argue that an ideal docMT metric should be (1) context-aware—capturing document-level coherence and accuracy, (2) structured—evaluating aspects such as fluency, accuracy, and coherence separately, and (3) interpretable—explicitly identifying translation errors for clear, objective human evaluation. To this end, we design a GPT-4-based evaluation pipeline to provide deeper insights into the docMT capabilities of LLMs.
@@ 2 Problem Settings
  • DOC: We instruct the LLM to directly translate the entire documented in one pass.
The DOC approach is designed to capture inter-sentence dependencies by considering the full document context, potentially leading to more coherent and accurate translations. However, it also requires understanding and generating longer text tokens, which increases the risk of cumulative errors if LLMs are not specialized into document-level translation yet.
@@ Results.
Comparison Cases in ST3 and DOC Prediction in ST3: The 13th Provincial Tea Expo opened today… This morning, the 13th Hunan Tea Industry Expo and… were held in the Hunan International Convention… At the opening ceremony, the provincial leaders awarded the fourth batch of enterprises with the right to use the Lake South Red Tea trademark. New Hope held the Lake Red Source. The Chaozhou tea industry warmly welcomed the arrival of the new spring.   Prediction in DOC: The 13th Hunan Tea Industry Expo opened today at the Hunan International Convention… At the opening ceremony, Hunan’s provincial leaders awarded the fourth batch of enterprises with the right to use the Hunan Red Tea trademark… The Hunan Red Tea is red in color and has injected the cultural connotation of tea into it, making it popular and lively, and the Hunan tea industry is looking forward to a new spring.   Prediction in ST3: Color: As shown in the picture (please avoid shooting to avoid color difference Therefore, girls who can’t drive should not complain about their clothes being old…{CJK*} UTF8gbsn 2021.6.11部分圈中售出。看好编号下单,古董物品售出不退不换。购买须知The products sold at this store are non-refundable… The store does not accept styles that are different from what is imagined, and size and style cannot be used as reasons for refunds or exchanges…   Prediction in DOC: Color: As shown in the picture (Please note that the color difference may not be avoided due to shooting… So, some girls who can’t drive vintage clothing should not say that the clothes are old-fashioned, but that you are not suitable for it!… If the item is not suitable for personal reasons, such as not fitting or not liking it, you can ask the store owner to transfer it to the shelf, and once it is sold, it cannot be exchanged or refunded. Part of the circle in the middle was sold on June 11, 2021

### v3

# Fine-Grained and Multi-Dimensional Metrics for Document-Level Machine Translation

@@ Abstract
Large language models (LLMs) have excelled in various NLP tasks, including machine translation (MT), yet most studies focus on sentence-level translation. This work investigates the inherent capability of instruction-tuned LLMs for document-level translation (docMT). Unlike prior approaches that require specialized techniques, we evaluate LLMs by directly prompting them to translate entire documents in a single pass. Our results show that this method improves translation quality compared to translating sentences separately, even without document-level fine-tuning. However, this advantage is not reflected in BLEU scores, which often favor sentence-based translations. We propose using the LLM-as-a-judge paradigm for evaluation, where GPT-4 is used to assess document coherence, accuracy, and fluency in a more nuanced way than n-gram-based metrics. Overall, our work demonstrates that instruction-tuned LLMs can effectively leverage document context for translation. However, we caution against using BLEU scores for evaluating docMT, as they often provide misleading outcomes, failing to capture the quality of document-level translation  [^1] Our code and the outputs from GPT4-as-a-judge are available at https://github.com/EIT-NLP/BLEUless_DocMT.
Fine-Grained and Multi-Dimensional Metrics for Document-Level Machine Translation
@@ 1 Introduction
A key challenge in our research is the evaluation of document-level machine translation (docMT). Traditional metrics  [^2] While COMET Rei et al. (2020) is more reliable than BLEU for sentence-level translation, it is trained exclusively on sentence-level data. As a result, using COMET to evaluate docMT can be unreliable, since out-of-distribution. like BLEU  [^3] Although we do not want to use BLEU based metric, it remains a common metric in existing/recent research, despite its limitations., ChrF, and TER Papineni et al. (2002); Popović (2015); Snover et al. (2006), though widely used, often poorly correlate with human judgment Freitag et al. (2022), especially in docMT, where maintaining coherence and logical flow across a document is essential—something n-gram overlap struggles to capture. Metrics like cTT, aZPT, and Blonde Jiang et al. (2021); Wang et al. (2023) address specific aspects such as terminology consistency and zero-pronoun accuracy, but still rely heavily on word matching and symbolic statistics. We argue that an ideal docMT metric should be (1) context-aware—capturing document-level coherence and accuracy, (2) structured—evaluating aspects such as fluency, accuracy, and coherence separately, and (3) interpretable—explicitly identifying translation errors for clear, objective human evaluation. To this end, we design a GPT-4-based evaluation pipeline to provide deeper insights into the docMT capabilities of LLMs.
@@ 2 Problem Settings
  • DOC: We instruct the LLM to directly translate the entire document in one pass.
The DOC approach is designed to capture inter-sentence dependencies by considering the full document context, potentially leading to more coherent and accurate translations. However, this approach requires the LLM to process and generate longer sequences of text, which can increase the risk of cumulative errors, especially if the model has not been explicitly optimized for document-level translation.
@@ Results.
Comparison Cases in ST3 and DOC Model: Vicuna-7B Prediction in ST3: The 13th Provincial Tea Expo opened today… This morning, the 13th Hunan Tea Industry Expo and… were held in the Hunan International Convention… At the opening ceremony, the provincial leaders awarded the fourth batch of enterprises with the right to use the Lake South Red Tea trademark. New Hope held the Lake Red Source. The Chaozhou tea industry warmly welcomed the arrival of the new spring.   Prediction in DOC: The 13th Hunan Tea Industry Expo opened today at the Hunan International Convention… At the opening ceremony, Hunan’s provincial leaders awarded the fourth batch of enterprises with the right to use the Hunan Red Tea trademark… The Hunan Red Tea is red in color and has injected the cultural connotation of tea into it, making it popular and lively, and the Hunan tea industry is looking forward to a new spring.   Model: Vicuna-7B-16K Prediction in ST3: Color: As shown in the picture (please avoid shooting to avoid color difference Therefore, girls who can’t drive should not complain about their clothes being old…{CJK*} UTF8gbsn 2021.6.11部分圈中售出。看好编号下单,古董物品售出不退不换。购买须知The products sold at this store are non-refundable… The store does not accept styles that are different from what is imagined, and size and style cannot be used as reasons for refunds or exchanges…   Prediction in DOC: Color: As shown in the picture (Please note that the color difference may not be avoided due to shooting… So, some girls who can’t drive vintage clothing should not say that the clothes are old-fashioned, but that you are not suitable for it!… If the item is not suitable for personal reasons, such as not fitting or not liking it, you can ask the store owner to transfer it to the shelf, and once it is sold, it cannot be exchanged or refunded. Part of the circle in the middle was sold on June 11, 2021